If 58.26 g of iodine reacts with excess aluminum. If 56.11 g of aluminum iodide is actually formed in the reaction percent yield of aluminum iodide
1
Expert's answer
2020-08-27T05:04:44-0400

The reaction is as following:

2Al + 3I2 = 2AlI3

The theoretical yield of the reaction is:

m(AlI3) = m(I2) × [2 × Mr(AlI3)] / [3 × Mr(I2)] = 58.26 g × 2 × 407.695 g/mol / [3 × 253.8089 g/mol] = 62.39 g

As experimental yield equals 56.11 g and theoretical yield is 62.39 g, percent yield of aluminium iodide equals:

w = (56.11 g / 62.39 g) × 100% = 89.93%


Answer: 89.93%